# Designing Speech Technologies for Australian Aboriginal English: Opportunities, Risks and Participation

### v2

@@ 7. Ethical considerations
Appendix B: Composition of corpus of Aboriginal English on YouTube
For readers who wish to hear examples of Aboriginal English, we recommend searching on YouTube using queries such as “My art, My culture”, “Telling Our Stories, Our Stolen Generations”, “Through our Eyes, Series 3”, or “Community Spotlight Artists Profile Cairns Indigenous arts fair”.
@@
Aboriginal English (hours)
@@
Table 3. Percent of speech in Aboriginal English in videos. Only videos with at least 50% were considered; 70+% was preferred.
@@
3. Psychological harms from ASR errors Primary goal of the project: reducing ASR errors for Aboriginal English. We also call for more HCI research both on potential harms experienced by Indigenous communities, as well as on potential mitigations in the application layer.
@@
Table 4. Project Risks and Mitigations
@@
Table 5. Project Opportunities and Enabling Factors

### v3

@@ Acknowledgements.
Appendix B: Composition of corpus of Australian Aboriginal English on YouTube
For readers who wish to hear examples of Australian Aboriginal English, we recommend searching on YouTube using queries such as “My art, My culture”, “Telling Our Stories, Our Stolen Generations”, “Through our Eyes, Series 3”, or “Community Spotlight Artists Profile Cairns Indigenous arts fair”.
@@
Australian Aboriginal English (hours)
@@
Table 3. Percent of speech in Australian Aboriginal English in videos. Only videos with at least 50% were considered; 70+% was preferred.
@@
3. Psychological harms from ASR errors Primary goal of the project: reducing ASR errors for Australian Aboriginal English. We also call for more HCI research both on potential harms experienced by Indigenous communities, as well as on potential mitigations in the application layer.
@@
Table 4. Project risks and mitigations.
@@
Table 5. Project opportunities and enabling factors.